OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. Please Help! 16x^2 + 8x + 1 = 0

OpenStudy (anonymous):

let "d" be the discriminant d=8^2-4*16=0 so the equation has double root x=(-8)/(2*16)=-8/32=-1/4 so 16x^2+8x+1=16(x+1/4)^2=0 i.e. x=-1/4

OpenStudy (anonymous):

I still don't get it.. can you explain it to me again?

OpenStudy (anonymous):

to calculate the discriminant "d" of an equation ax^2+bx+c=0 d=b^2-4*a*c in our case a=16,b=8 and c=1 so d=8^2-4*16=64-64=0 is it clear till now ?

OpenStudy (anonymous):

Okey, I'm starting to understand it now, thanks.